The team got together today to work on various open tasks and to experiment with some of the applications on the system.
We cleaned one of the mechanical mice loaned to us by Xerox PARC and it works quite well. We cleaned a second similar mouse but a bearing near the top of the ball cavity failed making it perform erratically. Several of the balls in the bearing fell out - each barely visible to the naked eye - and it does not look like it can be repaired. Finally, we still need to acquire and connect a DE19 plug to the optical mouse that came with the system.
We serviced the disk drive, replacing the air filter and adding a touch of the oil to the wipers on the arm bearings. The drive is working quite well.
We undertook some scoping and capture of the ethernet connection, to provide final specs for Ken's ethernet adapter project. We weren't seeing what we expected, so the team hooked up the logic analyzer to probe the adapter more fully.
ALTO DISK TOOL
I am away on holiday for the next week, having to put aside the work while I pack tonight.